Solve equations by factoring
Principle of zero products
Application problems
Number Sense: The sum of two number is six. The sum of the squares of the two numbers is twenty. Find the two numbers.
a + b = 6
a^2 + b^2 = 20
Don't know if I set it up right and what to do next.

Answer by lwsshak3(11628)

Number Sense: The sum of two number is six. The sum of the squares of the two numbers is twenty. Find the two numbers.
a + b = 6
a^2 + b^2 = 20
**
You set up correctly. I will just continue where you left off.
a+b=6
a=(6-b)
sub(6-b) for a in second equation
(6-b)^2 + b^2 = 20
36-12b+b^2+b^2=20
2b^2-12b+16=0
b^2-6b+8=0
(b-4)(b-2)=0
b=4
a=6-b=2
or b=2
a=6-b=4
curves intersect at (4,2) and (2,4)